Hi,
Bill sold 2 used books for $452 each. However the cost of each book was different. As a result a profit of 22% was made on the first book sold but a loss of 22% was made on the second sale. What is Bill's total profit or loss on both books.
Can you please explain how to do this problem?

Answer by scott8148(6628)

let f="cost of 1st book" and s="cost of 2nd book"

the profit (or loss) on an item is the difference between the cost and the selling price

the percentage profit/loss is the amount of profit/loss divided by the cost

.22 = (452-f) / f
multiplying by f ___ .22f = 452-f
adding f ___ 1.22f = 452
dividing by 1.22 ___ f = 370.49
profit = 452 - 370.49 = 81.51

-.22 = (452-s) / s
multiplying by s ___ -.22s = 452-s
adding s ___ .78s = 452
dividing by .78 ___ s = 579.49
loss = 452 - 579.49 = -127.49

so, the total LOSS on both books is 81.51 - 127.49 or -45.98